I,Λ2, X3 , X4,X5 be a sample from N
Answer #1

Here,

~ Normal(0, 1), 2-1, 2, 3, 4, 5

That implies,

sum_{i=1}^{5}rac{X_i^2 }{4} sim chi^2_{1}, i = 1,2,3,4,5

and by additive property of chi square distribution,

sum_{i=1}^{5}rac{X_i^2 }{4} sim chi^2_{5}

To find

5 Pl 〉 X > 5,75)

5 = 1-P(〉 X < 5.75)

= 1 - P(sum_{i=1}^{5}rac{X_i^2}{4} leq rac{5.75}{4})

= 1 - P(chi^2_{5} leq 1.43750)

= 1 - 0.0798324

= 0.920168